OpenStudy (anonymous):

What is the exact value of the expression 6(sqrt y^8) − (sqrt9y^4) + 3(sqrt y^8) −(sqrty^4) ? Simplify if possible.

OpenStudy (anonymous):

\[6\sqrt{y ^{8}}-\sqrt{9y ^{4}}+3\sqrt{y ^{8}}-\sqrt{y ^{4}}\]

jhonyy9 (jhonyy9):

sqrty8 =y4 sqrt9y4=3y2 sqrty4=y2

OpenStudy (anonymous):

thank you :)

jhonyy9 (jhonyy9):

6y4 -3y2 +3y4

jhonyy9 (jhonyy9):

-y2

OpenStudy (anonymous):

no, it is positive :) u were right the 1st time

jhonyy9 (jhonyy9):

=9y4-4y2

jhonyy9 (jhonyy9):

yes this -y2 is continue

jhonyy9 (jhonyy9):

of last line

OpenStudy (anonymous):

What is the product of \[2\sqrt{36d}*3d \sqrt{2}\]? :)

jhonyy9 (jhonyy9):

12sqrtd *3dsqrt2

jhonyy9 (jhonyy9):

=36dsqrt2d